A certain force accelerates a 4.2kg mass by 5m/s2. What would be the acceleration of 3.4kg mass when the same forces acts on it?
1
Expert's answer
2020-08-31T12:50:16-0400

"a_1 = 5 m\/s^2"

"m_1 = 4.2 kg"

"m_2 = 3.4 kg"

The equation for acceleration:

"a = \frac{F}{m}"

"F = a\times m"

"F = 5\times 4.2 = 21 N"

"a_2 = \frac{21 N}{3.4 kg} = 6.17 m/s^2"

Answer: 6.17 m/s2